# reference.bib

### v2

@@ I INTRODUCTION
  • Experiments are conducted on various benchmarks, including ETH/UCY and Stanford Drone Dataset, and the results attained demonstrate the state-of-the-art performance of our proposed framework.
@@ III-D Training and Loss Function
Our training scheme consists of two stages to train two modules separately. The goal estimation module is trained using Binary Cross-Entropy loss, which measures the dissimilarity between the ground truth of future trajectory heat-map $\mathcal{H}_{f}$ and predicted heat-map ~$\widetilde{\mathcal{H}}_{f}$:
@@
After the goal estimation module is trained, we freeze it and train the trajectory prediction module. The training objective of the trajectory prediction module is to learn a distribution $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ for each step $k$ of diffusion, following the regular setting of diffusion models [DDPM_2020_Neurips, DDIM_2021_ICLR]:
@@
TABLE I: Quantitative results on Stanford Drone Dataset (SDD).
Method ADE20 FDE20
Goal-GAN [Goal-GAN_2020_ACCV] 12.20 22.10
MG-GAN [MG-GAN_2021_ICCV] 13.60 25.80
PECNet [PECNet_2020_ECCV] 9.96 15.88
Y-net [YNet_2021_ICCV] 7.85 11.85
Goal-SAR [GoalSAR_2022_CVPRW] 7.75 11.83
MID [MID_2022_CVPR] 7.61 14.30
LED [LED_2023_CVPR] 8.48 11.66
GDTS (Ours) 7.39 11.80
TABLE II: Quantitative results on ETH/UCY dataset. †: The results after the data leakage issue is fixed.
@@
TABLE III: Ablation study on sampling algorithm of diffusion and $K_{t}$. The last row is the setting of our final reported results.
@@
DDPM - 7.70 11.98 $\sim$65
DDIM - 7.81 12.13 $\sim$24
TS 5 7.56 11.94 $\sim$24
TS 50 7.52 12.15 $\sim$21
TS 20 7.39 11.80 $\sim$24
IV RESULTS
IV-A Experiments and Datasets
To evaluate the performance of our model, we conducted experiments on two real-world pedestrian datasets: the ETH/UCY dataset and the Stanford Drone Dataset (SDD). These datasets capture the movement of pedestrians from a bird’s-eye view perspective and positions are manually annotated. The sample rate of datasets is 2.5 FPS. Given the pedestrian position of the last $t_{h}=8$ frames, the task is to predict the trajectory of the following $t_{f}=12$ frames.
Datasets The ETH/UCY [ETH_2009_ICCV, UCY_2007] dataset consists of 5 scenes: ETH, Hotel, Univ, Zara1, and Zara2. Following the common leave-one-scene-out strategy in previous work [SocialGAN_2018_CVPR, YNet_2021_ICCV], we use four scenes for training and the remaining one for testing. Stanford Drone Dataset [SDD_2016_ECCV] is a large-scale dataset that contains 11,216 unique pedestrians on the university campus. We use the same dataset split as several recent works [GoalSAR_2022_CVPRW, PECNet_2020_ECCV]: 30 scenes are used for training, and the remaining 17 scenes are used for testing.
Implementation Details All the experiments were conducted on an NVIDIA 3080Ti GPU with PyTorch implementation. The goal predictor is implemented using a 5-layer U-Net as the backbone. The architecture of the diffusion network is a three-layer transformer encoder similar to MID [MID_2022_CVPR], and readers can refer to the original paper for more details. We adopt a batch size of $64$. An Adam optimizer is employed with an initial learning rate of $10^{-3}$, and exponential annealing is applied for the learning rate. The goal estimation module is trained for $150$ epochs and the trajectory prediction module is trained for $250$ epochs. The DDPM diffusion step $K$ is $100$, the DDIM diffusion step $K_{I}$ is $20$, and the trunk step $K_{t}$ is $20$. Additionally, we utilize Test-Time-Sampling-Trick (TTST) in our model to improve the accuracy of goal estimation [YNet_2021_ICCV], and the number of TTST samples $N_{ttst}$ is selected to be $1000$. The final number of predictions $N=20$.
IV-B Quantitative Results
Baseline We compare GDTS to different methods: Goal-based methods include Goal-GAN [Goal-GAN_2020_ACCV], MG-GAN [MG-GAN_2021_ICCV], PECNet [PECNet_2020_ECCV], Y-net [YNet_2021_ICCV], Goal-SAR [GoalSAR_2022_CVPRW]. We do not include NSP-SFM [NSPSFM_2022_ECCV] here due to its unfair hyperparameters tuning for each scene in ETH/UCY. MID [MID_2022_CVPR], LED [LED_2023_CVPR] and SingularTrajectory [Singulartrajectory_CVPR_2024] are methods which leverage diffusion model for pedestrian trajectory prediction. Note that SingularTrajectory [Singulartrajectory_CVPR_2024] only reports official results in ETH/UCY.
Discussion Results of experiments on the SDD are reported in Table I. With the explicit leveraging of goal information, the FDE20 of GDTS is only larger than LED by a small margin, and smaller than all of the other compared methods. Meanwhile, our method achieves the best ADE20, further confirming the effectiveness of GDTS. Furthermore, the results in Table II show that our model achieves comparable accuracy in the ETH/UCY dataset. Since SDD is substantially larger than the ETH/UCY dataset, we argue that our method is more suitable for datasets with larger sizes.
TABLE IV: Ablation study on combination scheme of tree sampling algorithm. The last row is the setting of our final reported results.
@@ IV-B Quantitative Results
d-DDPM DDPM 20 80 7.70 11.95 $\sim$65
d-DDPM d-DDPM 20 80 7.36 11.63 $\sim$65
DDIM DDIM 4 16 8.02 12.62 $\sim$21
d-DDPM DDIM 20 16 7.39 11.84 $\sim$24
@@ IV-C Ablation Study
Sampling Algorithm of Diffusion To investigate the influence of different sampling algorithms, we replace tree sampling (TS) with DDPM and DDIM, and conduct the experiment on SDD. We report the average inference time for one agent in addition to the ADE20 and FDE20. Since LED [LED_2023_CVPR] and SingularTrajectory [Singulartrajectory_CVPR_2024] do not have an official implementation nor reported inference time on SDD, we only include MID which uses DDPM for comparison. The results in Table III verify that our proposed tree sampling algorithm performs the best and has a shorter inference time than DDPM and DDIM, indicating the superiority of our sampling algorithm. Compared with MID, our GDTS method reduces the inference time from 139 ms to around 25 ms.
Trunk Step of Tree Sampling We also explore the effect of the trunk step $K_{t}$. As shown in Table III, $K_{t}$ does not significantly influence the inference time since we use DDIM in the branch stage which already has a fast sampling speed. However, the accuracy of our method decreases when $K_{t}$ is either too large or too small. When $K_{t}$ is large, small $K_{b}$ makes it challenging for the branch stage to drive the roughly denoised trajectory to various modalities, limiting the diversity of the results. When $K_{t}$ is small, the branch stage that uses DDIM undertakes most of the denoising task, without obtaining enough guidance from the trunk stage, which degrades the performance.
Combination scheme of Tree Sampling We compare different combinations of DDPM, d-DDPM, and DDIM in tree sampling algorithm to verify the effectiveness of our scheme. Note that the result of the trunk stage must be deterministic, hence only d-DDPM and DDIM ($\eta=0$) can be used in the trunk stage. Results in Table IV demonstrate that using d-DDPM in both stages gets the lowest ADE20 and FDE20, probably because d-DDPM focuses on reconstruction rather than generation, i.e., accuracy rather than diversity. However, the inference time is much longer than our scheme, and our scheme achieves a good balance between prediction accuracy and inference speed.
@@ IV-D Qualitative Results
Fig. 3 presents some qualitative results of our GDTS method on the ETH/UCY and SDD datasets. The target pedestrian in the scene (a) turns right while moving straight in the scene (b) and (c). From the left column, it can be observed that the goals from the final prediction (in cyan) are closer to the ground truth (in yellow) compared to the goal estimations obtained from the goal estimation module (in blue), indicating that the diffusion models can improve FDE. The middle column shows the common goal $g_{*}$ and the roughly denoised trajectory generated by the trunk stage, demonstrating that the common goal can guide the denoised trajectory growing toward the correct direction. The right column shows the final prediction results, demonstrating the ability of GDTS to generate future trajectory predictions with different modalities.
@@ V CONCLUSIONS
In conclusion, GDTS is introduced in this work for multi-modal pedestrian trajectory prediction. The proposed development is a framework that integrates goal estimation and the diffusion probabilistic model. Our novel tree sampling algorithm leverages common feature for multiple modalities generation to accelerate inference speed. Experiments on real-world datasets demonstrate that our GDTS method can predict multiple scene-compliant trajectories and achieve state-of-the-art performance with real-time inference speed in real-world datasets.
Limitation and Future Work The interaction between agents is currently not considered in this work, yet it is crucial for human motion prediction. Recent works [GoalSAR_2022_CVPRW, NSPSFM_2022_ECCV, PPT_ECCV_2024] indicate that recurrent prediction method could improve prediction accuracy, and we will explore how to integrate it into our framework in the future.

### v3

@@ I INTRODUCTION
  • We conduct a series of experiments on various datasets, including ETH/UCY, Stanford Drone Dataset and intersection Drone Dataset, and the results attained demonstrate the state-of-the-art performance of our proposed framework in large-size datasets.
@@ III-D Training and Loss Function
Our training scheme consists of two stages to train two modules separately. First, the goal estimation module is trained solely using Binary Cross-Entropy loss, which measures the dissimilarity between the ground truth of future trajectory heat-map $\mathcal{H}_{f}$ and predicted heat-map ~$\widetilde{\mathcal{H}}_{f}$:
@@
$\mathcal{L}_{1}=\mathcal{L}_{goal}$ (2)
In the second stage, we train the goal estimation module together with the trajectory prediction module. The training objective of the trajectory prediction module is to learn a distribution $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ for each step $k$ of diffusion, following the regular setting of diffusion models [DDPM_2020_Neurips, DDIM_2021_ICLR]:
@@
$\mathcal{L}_{2}=\lambda\mathcal{L}_{goal}+\mathcal{L}_{traj}$ (4)
IV RESULTS
IV-A Experiments and Datasets
To evaluate the performance of our model, we conducted experiments on three public real-world pedestrian datasets: the ETH/UCY dataset, the Stanford Drone Dataset (SDD) and the intersection Drone Dataset (inD). These datasets capture the movement of pedestrians from a bird’s-eye view perspective and positions are manually annotated. The sample rate of datasets is 2.5 FPS. Given the pedestrian position of the last $t_{h}=8$ frames, the task is to predict the trajectory of the following $t_{f}=12$ frames.
TABLE I: Quantitative results on ETH/UCY dataset. †indicates the model after the data leakage issue is fixed.
@@ IV-A Experiments and Datasets
TABLE II: Quantitative results on Stanford Drone Dataset (SDD). Bold number indicates the best.
Method ADE20 FDE20
Goal-GAN [Goal-GAN_2020_ACCV] 12.20 22.10
MG-GAN [MG-GAN_2021_ICCV] 13.60 25.80
PECNet [PECNet_2020_ECCV] 9.96 15.88
Y-net [YNet_2021_ICCV] 7.85 11.85
Goal-SAR [GoalSAR_2022_CVPRW] 7.75 11.83
MID [MID_2022_CVPR] 7.61 14.30
LED [LED_2023_CVPR] 8.48 11.66
GDTS (Ours) 7.42 11.57
Datasets The ETH/UCY [ETH_2009_ICCV, UCY_2007] dataset consists of 5 scenes: ETH, Hotel, Univ, Zara1, and Zara2. Following the common leave-one-scene-out strategy in previous work [SocialGAN_2018_CVPR, YNet_2021_ICCV], we use four scenes for training and the remaining one for testing. SDD [SDD_2016_ECCV] is a large-scale dataset that contains 11,216 unique pedestrians on the university campus. We use the same dataset split as several recent works [GoalSAR_2022_CVPRW, PECNet_2020_ECCV]: 30 scenes are used for training, and the remaining 17 scenes are used for testing. inD [ind_IV_2020] includes 4 different intersection scenes in German and we follow the same evaluation protocol applied in  [GoalSAR_2022_CVPRW, ACVRNN_2021, gsgformer_2023gsgformer], where only pedestrian trajectories are retained and split into train, validation, and test sets by a 70%/10%/20% portion.
Implementation Details All the experiments were conducted on an NVIDIA 3080Ti GPU with PyTorch implementation. The goal predictor is implemented using a 5-layer U-Net as the backbone. The architecture of the diffusion network is a three-layer transformer encoder similar to MID [MID_2022_CVPR], and readers can refer to the original paper for more details. We adopt a batch size of $64$. An Adam optimizer is employed with an initial learning rate of $10^{-3}$, and exponential annealing is applied for the learning rate. The goal estimation module is trained for $150$ epochs first and then trained with trajectory prediction module for $250$ epochs. The DDPM diffusion step $K$ is $100$, the DDIM diffusion step $K_{I}$ is $20$, and the trunk step $K_{t}$ is $30$. Additionally, we utilize Test-Time-Sampling-Trick (TTST) in our model to improve the accuracy of goal estimation [YNet_2021_ICCV], and the number of TTST samples $N_{ttst}$ is selected to be $1000$. The goal estimation loss coefficient $\lambda=20$ and The final number of predictions $N=20$.
TABLE III: Quantitative results on Intersection Drone Dataset (inD). Bold number indicates the best. * indicates the result reported in [GoalSAR_2022_CVPRW]
Method ADE20 FDE20
Social-GAN [SocialGAN_2018_CVPR] 0.48 0.99
ST-GAT [STGAT_2019_ICCV] 0.48 1.00
AC-VRNN [ACVRNN_2021] 0.42 0.80
GSGFormer [gsgformer_2023gsgformer] 0.30 0.55
Y-net* [YNet_2021_ICCV] 0.34 0.56
Goal-SAR [GoalSAR_2022_CVPRW] 0.31 0.54
GDTS (Ours) 0.30 0.53
IV-B Quantitative Results
Baseline We compare GDTS to different methods: Goal-based methods include Goal-GAN [Goal-GAN_2020_ACCV], MG-GAN [MG-GAN_2021_ICCV], PECNet [PECNet_2020_ECCV], Y-net [YNet_2021_ICCV], Goal-SAR [GoalSAR_2022_CVPRW]. We do not include NSP-SFM [NSPSFM_2022_ECCV] here due to its unfair hyperparameters tuning for each scene in ETH/UCY. MID [MID_2022_CVPR], LED [LED_2023_CVPR] and SingularTrajectory [Singulartrajectory_CVPR_2024] are methods which leverage diffusion model for pedestrian trajectory prediction. Specifically for inD, besides goal-based methods, we also compare our model with some SOTA methods: Social-GAN [SocialGAN_2018_CVPR], ST-GAT [STGAT_2019_ICCV], AC-VRNN [ACVRNN_2021] and GSGFormer [gsgformer_2023gsgformer], which are all training with the same dataset split as our model. Note that recent works with different evaluation protocal [hierarchical_TVT_2024] are not included for fair comparison.
TABLE IV: Ablation study on sampling algorithm of diffusion and $K_{t}$. The last row is the setting of our final reported results. Bold number indicates the best.
@@ IV-B Quantitative Results
DDPM - 7.76 11.91 $\sim$65
DDIM - 7.91 12.01 $\sim$24
TS 5 7.61 11.88 $\sim$24
TS 50 7.56 12.08 $\sim$21
TS (Ours) 30 7.42 11.57 $\sim$24
Discussion Results in Table I show that our model achieves comparable accuracy in the ETH/UCY dataset. Furthermore, results of experiments on the SDD are reported in Table II. With the combination of explicit guide guidance and diffusion denoising, our model achieves the best ADE20 and FDE20 among all compared methods. Results in inD shown in Table III are in line with the previous table, further confirming the effectiveness of GDTS. Since SDD and inD are substantially larger than the ETH/UCY dataset, we argue that our method is more suitable for datasets with larger sizes.
TABLE V: Ablation study on combination scheme of tree sampling algorithm. The last row is the setting of our final reported results. Bold number indicates the best.
@@
d-DDPM DDPM 30 70 7.74 11.83 $\sim$65
d-DDPM d-DDPM 30 70 7.36 11.63 $\sim$65
DDIM DDIM 6 14 8.02 12.62 $\sim$21
d-DDPM DDIM 30 14 7.42 11.57 $\sim$24
@@ IV-C Ablation Study
Sampling Algorithm of Diffusion To investigate the influence of different sampling algorithms, we replace tree sampling (TS) with DDPM and DDIM, and conduct the experiment on SDD. We report the average inference time for one agent in addition to the ADE20 and FDE20. Since LED [LED_2023_CVPR] and SingularTrajectory [Singulartrajectory_CVPR_2024] do not have an official implementation nor reported inference time on SDD, we only include MID which uses DDPM for comparison. The results in Table IV verify that our proposed tree sampling algorithm performs the best and has a shorter inference time than DDPM and DDIM, indicating the superiority of our sampling algorithm. Compared with MID, our GDTS method reduces the inference time from 139 ms to around 25 ms.
Trunk Step of Tree Sampling We also explore the effect of the trunk step $K_{t}$. As shown in Table IV, $K_{t}$ does not significantly influence the inference time since we use DDIM in the branch stage which already has a fast sampling speed. However, the accuracy of our method decreases when $K_{t}$ is either too large or too small. When $K_{t}$ is large, small $K_{b}$ makes it challenging for the branch stage to drive the roughly denoised trajectory to various modalities, limiting the diversity of the results. When $K_{t}$ is small, the branch stage that uses DDIM undertakes most of the denoising task, without obtaining enough guidance from the trunk stage, which degrades the performance.
Combination scheme of Tree Sampling We compare different combinations of DDPM, d-DDPM, and DDIM in tree sampling algorithm to verify the effectiveness of our scheme. Note that the result of the trunk stage must be deterministic, hence only d-DDPM and DDIM ($\eta=0$) can be used in the trunk stage. Results in Table V demonstrate that using d-DDPM in both stages gets the lowest ADE20 and FDE20, probably because d-DDPM focuses on reconstruction rather than generation, i.e., accuracy rather than diversity. However, the inference time is much longer than our scheme, and our scheme achieves a good balance between prediction accuracy and inference speed.
@@ IV-D Qualitative Results
Fig. 3 presents the qualitative results of our GDTS method on the ETH/UCY and SDD datasets. The target pedestrian in the scene (a) turns right while moving straight in the scene (b) and (c). From the left column, it can be observed that the goals from the final prediction (in cyan) are closer to the ground truth (in yellow) compared to the goal estimations obtained from the goal estimation module (in blue), indicating that the diffusion models can improve FDE. The middle column shows the common goal $g_{*}$ and the roughly denoised trajectory generated by the trunk stage, demonstrating that the common goal can guide the denoised trajectory growing toward the correct direction. The right column shows the final prediction results, demonstrating the ability of GDTS to generate future trajectory predictions with different modalities.
@@ V CONCLUSIONS
In conclusion, GDTS is introduced in this work for multi-modal pedestrian trajectory prediction. The proposed development is a framework that integrates goal estimation and the diffusion probabilistic model. Our novel tree sampling algorithm leverages common feature for multiple modalities generation to accelerate inference speed. Experiments on real-world datasets demonstrate that our GDTS method can predict multiple scene-compliant trajectories and achieve state-of-the-art performance with real-time inference speed in real-world datasets. Nevertheless, the interaction between agents is currently not considered in this work, yet it is crucial for human motion prediction. Predicting additional probabilistic score for each modality could help with downstream decision-making and planning. In this sense, we will explore the integration of multi-agent interactions into our framework.